OpenStudy (anonymous):

Write a system of two equations using x and y in each for the following. The sum of two numbers is 170 and their difference is 26.

OpenStudy (anonymous):

x+y=170 x-y=26

OpenStudy (anonymous):

Let the two numbers be x and y. x + y is given as 170, so we put x + 7 = 170. The difference is 26, so we put x - 7 = 26 That is the system of two equations.

OpenStudy (anonymous):

ok typos: Let the two numbers be x and y. x + y is given as 170, so we put x + y = 170. The difference is 26, so we put x - y = 26 That is the system of two equations.

OpenStudy (anonymous):

Note, that we could also frame the system as: y + x = 170 y - x = 26 Hope I am not confusing you ..

OpenStudy (anonymous):

Hey you from india sneha

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ME too names are so recogniseale would you join me on facebook

OpenStudy (anonymous):

thanks both of you:)